Gail can type a chapter book in 4 hours jack can type a chapter book in 10 hours. How long would it take them to type a chapter book together?
Answer by stanbon(75887)

Gail can type a chapter book in 4 hours jack can type a chapter book in 10 hours. How long would it take them to type a chapter book together?
------------
Gail DATA:
time = 4 hrs/job ; rate = 1/4 job/hr
----------------------------------
Jack DATA:
time = 10 hrs/job ;rate = 1/10 job/hr
-----------------------------------
Together DATA:
time = x hrs/job ; rate = 1/x job/hr
-----------------------------------
EQUATION:
rate + rate = together rate
1/4 + 1/10 = 1/x
5x + 2x = 20
7x = 20
x = 20/7 = 2.857.. hrs (time to do the job together)
============
Cheers,
Stan H.